. A compound with a molar mass of about 28 g/mol contains 85.7% carbon and 14.3% hydrogen by mass. Write the Lewis structure for a molecule of the compound.
Answer #1

Molay mass of covepound = 287/mol In this c-85.7% C molar mass = 129 m. tersg 100% ~ 289 85.9% = 3 (289) (85.77.) = 23.996 =Ĉ Valence electrons = 4 # Valence elections = 1 Total Valence electrons = 4 + 1(H) = 8 Lewis structure çi : A IU-I H CScanned